XCO DE Problems 1-3 are based on the following description Electrical energy is applied to a system enclosed inside the pistoQuestion 2 The expansion work in J is closer to: 20 30 O 300 O 0.2 2000 4000 400 200 Question 3Question 3 The increase in energy of the system in J is closer to: 90 200 110 600 3200 O 400 60 33
Answer #1

Given: I = 2 amps ; V= 10 volts; Time(t) =20 sec ; P= 200 KPa= constant; dv= 0.2m^{3} ;Q= -90 J

(1)The amount of electric energy applied to the system is given by:

W_{elec}=V*I*t=10*2*20=400J

(2)The expansion work is:

W_{exp}=Pdv=200*1000*0.2=40000 J

(3)The increase in energy of system(dE):

From 1st law of thermodynamics:

dQ=dE+dW

Q=dE+W_{exp}-W_{elec} ( since the electrical energy is applied to the system so it is taken as negative and the expansion is done by the system so it is taken as positive)

=> -90=dE+40000-400

so, dE= -39690 J

Note: ques (2) and (3) option doesn't match because something wrong with the data given in ques.